Reports 1-1 of 1 Clear search Modify search
AdV-SAT (Suspension control upgrade)
Boschi, Gennai, Piendibene - 16:03 Thursday 27 November 2025 (68266) Print this report
BS DSP Board Replacements

Investigating the glitches and the DAQ transmission problems of Sc_BS, using the RapidIO diagnostics data provided by its MCH, two DSP boards generating many errors were identified. One, connected to PSDi (serial #42084), was replaced yesterday morning, with a spare (serial #45261). The other, connected to the mirror, was a P5 board (serial #51181). This latter was replaced with a standard p4 board (serial #45257) with its output stage modified in order to have the same output resistance of a p5. In this way the gain was kept while the dynamics was reduced by only about 12 % (10 V / 28.5 Ohm = 0.35 mA vs 0.31 mA of maximum current erogation of AD8397 opamp). This morning the interferometer was successfully locked with this new setup.

Comments to this report:
rolland - 10:36 Friday 28 November 2025 (68276) Print this report

SInce the BS board replacement, the DMS flag monitoring the TF from Sc_BS_MIR_Z_CORR to the coil current monitoring channels (V1:Sc_BS_MIR_VOUT_DL and DR, UL, UR) is red.

Indeed, the VIM plots show that there is no coherence anymore between them at 2012 Hz (figure 1, top right).

Figure 2 : looking at the channels with dataDisplay:

- purple: current situtation (ITF locked)

- blue: situation on 24 November

--> It looks like we now see only ADC noise, and that the four VOUT monitoring channels are not enabled/plugged?

 

Images attached to this comment
Search Help
×

Warning

Error

The present report has been modified outside this window. Please check for its integrity in the main page.

Refreshing this page will move this report into drafts.

×